Exercise 07.51 Algorithmic Question 3 of 3 Check My Work (2 remaining) A market research firm conducts telephone surveys with a 40% historical response rate. What is the probability that in a new sample of 400 telephone numbers at least 150 individuals will cooperate and respond to the questions? In other words, what is the probability that the sample proportion will be at least 150/400 375 (to 4 decimals)? Use z-table.
Answer #1

Ans:

Sampling distribution of sample proportion:

mean=p=0.4

standard deviation=sqrt(0.4*(1-0.4)/400)=0.0245

sample proportion=150/400=0.375

z=(0.375-0.4)/0.0245=-1.021

P(z>=-1.021)=P(z<=1.021)=0.8464

*(without rounding off it will be 0.8463)